Use the following information to answer the next 2 questions. Jordan, a single woman, is very generous. She enjoys giving gifts to others and has given taxable gifts of $5,000,000 in prior years. This past year (2019) she gave the following gifts:

1. For Judy, she wrote a check in the amount of $48,000 to the local hospital to pay Judy's medical bills

2. To Mark, she gave a new car valued at $62,000

3. To Kristen, she gave a $19,000 painting

4. To Haley, she gave $35,000 to pay for Haley's tuition for the year at Louisiana State.

a. Calculate Jordan's total taxable gifts for the 2019 calendar year. Note that the Annual Exclusion Amount in 2019 was also $15,000 and that the Applicable Credit Amount in 2019 was $4,505,800.

A. $71,000

B. $104,000

C. $116,000

D. $121,000

E. $146,000

b. Calculate Jordan's gift tax liability due for the current year. Note that the Annual Exclusion Amount in 2019 was also $15,000 and that the Applicable Credit Amount in 2019 was $4,505,800

A. $0

B. $32,800

C. $18,900

D. $44,000

E. $58,400
Answer #1

Answer (a):

Correct answer is:

A. $71,000

Explanation:

1. For Judy, she wrote a check in the amount of $48,000 to the local hospital to pay Judy's medical bills: Qualified transfer

2. To Mark, she gave a new car valued at $62,000: Taxable amount = 62000 - 15000 = $47,000

3. To Kristen, she gave a $19,000 painting: Taxable amount = 19000 - 15000 = $4,000

4. To Haley, she gave $35,000 to pay for Haley's tuition for the year at Louisiana State : Taxable amount = 35000 - 15000 = $20,000

Taxable gift = 47000 + 4000 + 20000 = $71,000

Option A is correct and other options B, C and D are correct.

Answer (b):

Correct answer is:

A. $0

Explanation:

Jordan has applicable Credit Amount in 2019 was $4,505,800 and hence she can adjust Taxable gift of $71,000 against the same.

As such:

Jordan's gift tax liability due for the current year = $0

Option A is correct and other options B, C and D are correct.